# Antitrust, Amazon, and Algorithmic Auditing

### v1

@@ Introduction
From the perspective of individual users, digital platforms serve as both gateways to information and architects of their interaction spaces. Digital platforms wield considerable influence over what is commonly termed “choice architectures” (Thaler and Sunstein, 2008), the environments in which users navigate decision-making processes. Extensive research in psychology and related fields has demonstrated that the ability to manipulate choice architectures can profoundly shape user behavior and outcomes (Thaler and Sunstein, 2008). Digital platforms may use technical choice architectures expressed in algorithms to shape consumer preferences, search behavior and purchase patterns.
@@
That said, obtaining information about business models is usually not trivial. As a first step, a regulator or researcher seeking to assess the operations of a digital platform with the help of digital tools would typically try to access the platform through a regular web interface or engage in automated scraping of the platform’s content. However, interfaces of digital platforms have become more interactive, oftentimes rendering simple web access insufficient to capture the inner workings of a digital platform. Digital platforms and markets enabled by these platforms are also growing more complex, making it harder to gain a proper understanding of platforms’ business practices.
@@
In response, empirically oriented computer science research has developed various approaches to audit digital platforms over the last few years. Sandvig et al. (2014) distinguish between “code audits” (where researchers directly inspect a copy of the relevant algorithm), “noninvasive user audits” (where users interacting with the platform answer questions about what they did online), “scraping audits” (where researchers query a platform repeatedly, observing results), “sock puppet audits” (where researchers use computer programs to impersonate users), and “crowdsourced audits” (where researchers employ users as testers). This development points towards a future of algorithmic auditing, where potential adverse effects of algorithmic systems are investigated and evaluated using algorithmic auditing tools (Sandvig et al., 2014; Bandy, 2021; Metaxa et al., 2021). “The Answer to the Machine is in the Machine", as Charles Clark pronounced in the 1990’s (Clark, 1999): As digital platforms are increasingly relying of algorithmic systems to govern their digital platforms, regulators will have to resort increasingly to algorithmic systems to audit compliance on these platforms.
@@
This article explores some of the opportunities and challenges of algorithmic enforcement, using Amazon as an example. Recently, Amazon’s business policies and deployed algorithmic systems have come under intense scrutiny due to Amazon’s vertical integration, or in other words, its involvement in multiple stages of the production and distribution process for its products and services. In this article, we present a series of empirical investigations into the extent to which Amazon engages in practices that are typically described as “preferential treatment” or “self-preferencing.” In these investigations, we analyze different features of various choice architectures that Amazon presents to its customers. We investigate how the service decides the offer that is featured in its Buy Box (now called “Featured Offer”); the design feature of its page listing various offers for the same product; how it selects products offered to customers that use the conversational product search technology available through its virtual assistant technology Alexa; and how it displays results in its related-item recommendation system.
@@
Over the last years, Amazon has emerged as the largest online e-commerce platform by connecting millions of sellers across the world to billions of customers (Saqib, 2023). On the Amazon platform, the different stakeholders can be primarily classified into three categories: (a) sellers on the platform: they rely on Amazon to sell their products to consumers; (b) customers of the platform: they rely on Amazon for fulfilling their purchase needs; and (c) the platform organization itself: it mediates the interaction between the sellers and costumer.
@@
Second, in addition to the seller featured in the Amazon Buy Box, all other sellers are listed on a separate offer listing page (Dash et al., 2024). This page includes seller performance metrics such as average user ratings, percentage of positive feedback, or the number of ratings. This enables consumers to compare the sellers from whom they would like to purchase the product. We investigate the choice architecture of these offer listing pages in Section 4.2.
@@
While Amazon uses these and other choice architectures to design its platform, Amazon’s role is not restricted to a platform designer. In addition, Amazon sells directly to consumers in most of its marketplaces – for example, in the United States, France, or Germany – making it one of the major sellers on its own marketplace. This puts Amazon in direct competition with independent sellers for many products that are sold on Amazon. In some countries – for example, India – Amazon is not allowed to sell products on its marketplace directly. As a response, Amazon created joint ventures in India in collaboration with other retailers to sell products on Amazon.in. Cloudtail India and Appario private limited, for example, were two joint ventures which emerged as two of the biggest sellers on Amazon India. [^1] While Cloudtail ceased to offer products on Amazon in 2022, Appario remains in the business of selling products on Amazon. Both companies were sellers on Amazon India while the data presented in this article was collected.
@@
The intricate relationship between Amazon and sellers on Amazon Marketplace also plays out at the product level. Amazon manufactures its own private label products and sells them under different brand names (for example, AmazonBasics, Presto, or Solimo (Dash et al., 2021). Note that only Amazon sells Amazon’s private label products on Amazon. This policy is known as gating (Edelman, 2019).
In addition to Amazon’s role as a platform organizer, a seller, a fulfillment and logistical service provider, and a product manufacturer, Amazon finally maintains an advertising platform. When a customer looks for a product on Amazon, the platform presents organic and sponsored recommendations. While organic recommendations refer to the results that the algorithmic system produces in the desired context based on previous customer behavior, sponsored recommendations are sponsored product advertisements which are the result of second price auctions. On Amazon Marketplace, Amazon interleaves organic results with sponsored ones. If consumers were unable to sufficiently distinguish between both types of results, they might end up accidentally selecting advertised results, implicitly assuming that they were popular with other consumers. Moreover, as Amazon acts as both a product manufacturer and a seller on the platform, it may also utilize its own ad space to promote its own products. This arrangement could create incentives for Amazon to reserve particularly well-positioned recommendations for itself (Dash et al., 2021).
@@
The complex interactions between multiple players on Amazon are a result of continuous product development and innovation by Amazon over the last three decades. Originally, Amazon emerged as a book seller, entering the business of selling products via the Amazon platform itself only later. Similarly, major online travel agencies such as Booking.com or Expedia over time became vertically integrated with meta-search platforms, enabling these travel agencies to compete with independent competitors on these platforms (Cure et al., 2022).
@@
When a digital platform such as Amazon sells own products or products from affiliated producers on its online platform, this raises issues of vertical integration. Can a vertically integrated platform that operates both as platform provider and seller sufficiently distinguish between both roles? Or is there a risk that it has incentives to bias search results and recommendations in favor of its own or affiliated products, at the expense of products from competing independent sellers?
@@
Problems of vertical integration are not unique to digital platforms. The practice of supermarkets providing preferential shelf space to their own or affiliated products (see Tagliavini 2023, 202), and financial advisers who earn commissions offering potentially biased investment recommendations (Cookson et al., 2021; Egan, 2019), both raise similar questions. From an economic perspective, vertical integration can have various effects on social welfare. Vertical integration of digital platforms could increase consumer search costs, discourage product innovation and prompt competitors to leave the market, as preferential treatment by the platform provider for own or affiliated products may make it more difficult for independent competitors to compete (see Farronato, Fradkin, and MacKay 2023). However, it is not clear that such welfare-decreasing effects will always dominate. By engaging in preferential treatment, a dominant vertically integrated platform provider may have incentives to invest in product quality, product design and innovation. It may also have an incentive to decrease prices and increase output because it can capture a larger fraction of the benefits from such actions (often referred to as the “internalization of double mark-ups,” see European Commission 2008, 7). Therefore, whether vertically integrated platform providers are beneficial or detrimental to economic welfare is a complex question that is often impossible to answer in the abstract, based solely on theoretical considerations (see, for example, de Cornière and Taylor 2019; Hagiu, Teh, and Wright 2022; Anderson and Bedre-Defolie 2022).
@@
These insights from industrial organization research have had an influence on competition law’s treatment of vertical integration. With the increasing influence of the Chicago School in antitrust, many jurisdictions have become reluctant to condemn cases of vertical integration as such. Rather, antitrust authorities investigate potential violations of antitrust laws in vertical relationships as part of a rule-of-reason analysis. Under rule-of-reason analysis, antitrust agencies and courts examine both the positive and negative effects of a certain conduct on competition in order to determine whether the conduct violates antitrust laws. For example, U.S. courts have shifted from a per-se prohibition of resale price maintenance [^2] Compare Dr. Miles Medical Co. v. John D. Park & Sons Co., 220 U.S. 373 (1911) (treating a vertical price fixing scheme as per se illegal) with Leegin Creative Leather Prods., Inc. v. PSKS, Inc., 551 U.S. 877 (2007) (holding that all vertical distribution restraints – non-price and price – should be analyzed under the rule of reason). and territorial restrictions [^3] Compare United States v. Arnold, Schwinn & Co., 388 U.S. 365 (1967) (holding that vertical restraints limiting distributors to exclusive sales areas is per se illegal) with Cont’l T.V., Inc. v. GTE Sylvania, Inc., 433 U.S. 36 (1977) (applying a rule of reason analysis to non-price distribution restraints protecting interbrand competition). to a rule-of-reason analysis that focuses on the economic effects of vertical restraints. Relatedly, self-preferencing by a dominant firm is not a per se violation of European competition law (Art. 102 of the Treaty on the Functioning of the European Union, TFEU), but subjects to an effects test (Crémer, Montjoye, and Schweitzer, 2019, 7, 66).
@@
This effects-based approach to assessing vertical integration under antitrust laws has influenced the way the authorities have addressed alleged preferential treatment by digital platforms. The General Court of the European’s recent Google Shopping decision provides an example. In this decision, the court reviewed the European Commission’s finding that Google had promoted its own comparison shopping service on its general search engine at the expense of rival services (see General Court, Case T-612/17, Nov. 10, 2021, ECLI:EU:T:2021:763). In its decision, the court pointed out that not all foreclosure is detrimental to competition, and that the mere extension of a dominant position to a neighbouring market is not necessarily anti-competitive (General Court, Case T-612/17, Nov. 10, 2021, ECLI:EU:T:2021:763, para. 157, 162). While the court ultimately confirmed the European Commission’s decision, the decision demonstrates that whether self-preferencing violates European competition law can only be decided on a case-by-case basis. This analysis involves various different competition doctrines such as duty to deal, tying, margin squeeze, general abuse, or discrimination (see Colangelo 2023).
@@
Along these lines, the European Union introduced a regulatory system of digital markets that is based on an ex-ante approach where efficiency defenses raised by firms play a very limited role. With the Digital Markets Act (DMA), which came into full force in 2024, the European Union aims to make key digital markets fairer and more contestable by preventing large companies from abusing their market power and lowering barriers to the entry and expansion of smaller players. The DMA targets large digital platforms – so-called “gatekeepers” – such as Alphabet, Amazon, Apple, Meta, or Microsoft and imposes various obligations on them.
@@
Article 6(2) DMA prohibits digital platforms from using, in competition with other sellers, any non-public data that is generated or provided by those sellers. This prohibition is intended to countervail the ability of vertically integrated digital platforms to observe sales data from independent sellers with whom the platforms compete. The Digital Markets Act also includes prohibitions on combining data from different services belonging to the same company, provisions protecting advertisers and publishers using the gatekeeper against monopolization trends, provisions for ensuring interoperability and data access for business and individual users of platforms, and many others. If gatekeepers fail to abide by these obligations, the European Commission may impose fines of up to 10 % of the gatekeeper’s annual worldwide turnover (Art. 30(1) DMA). The Digital Market Act also empowers the Commission to monitor the effective compliance of digital platforms with the obligations of the DMA, potentially using independent external experts, auditors and member state authorities (Art. 26 DMA). Article 21 DMA entitles the European Commission to require access to data and algorithms of digital platforms.
Going beyond the Digital Markets Act, the European Union also enacted the Digital Services Act (DSA), which came into effect in 2024 as well. This regulation provides a comprehensive framework regarding illegal content, transparent advertising, and disinformation of online platform providers. Article 26 of the DSA, for example, requires online platforms to be transparent about the nature and origin of advertisements. This obligation also applies to “sponsored links” displayed by digital platforms. Articles 25 and 27 of the DSA include obligations on online interface design and recommender system transparency. The European Commission may impose fines of up to 6 % of the annual worldwide turnover of very large online providers (Art. 74(1) DSA), which includes Amazon, Alibaba, Apple, Google Shopping, and Zalando, among other providers. [^4] For a list of very large online platforms as designated by the European Commission, see https://digital-strategy.ec.europa.eu/en/policies/list-designated-vlops-and-vloses.
@@
We start our exploration of Amazon’s market practices by investigating how the service decides the offer that is featured in its Buy Box (now called “Featured Offer”). A typical Amazon product page has a rectangular box on the top-right corner, right next to the title of the product. This box is popularly known as the Buy Box. Figure 1 shows an example of such a Buy Box. The box contains the price of the product, a delivery estimate, and two buttons that read: ‘Add to Cart’ and ‘Buy Now’. The box also lists the name of a seller. In the Buy Box depicted in Figure 1, the seller is Appario Retail Private Ltd.
@@
Given the importance of the Buy Box for transactions on Amazon’s platform, Amazon exerts considerable power over customers’ purchase decisions through its algorithm that selects the seller that appears in the Buy Box. This decision is trivial if there is only one seller that offers a given product. However, in practice, multiple sellers can and often do offer the same product. In this scenario, Amazon selects the seller to be featured in the Buy Box through a proprietary algorithm known as the Featured Offer Selection algorithm, also commonly known as the Buy Box algorithm. The precise structure of this algorithm is unknown, making it a subject of speculation (and concern) among sellers on Amazon (Kaneshiro, 2020). Amazon presents some clues and recommendations on its seller interface regarding what factors sellers should consider to win the Buy Box. These factors include, for example, competitive pricing and seller performance metrics.
@@
To investigate this question, we first compare the price offered by Amazon and rival sellers in cases in which Amazon won the Buy Box. We find considerable numbers of instances in which Amazon won the Buy Box despite offering the product at a price that is higher than that offered by at least one rival seller: In the US and in India, 25% of the Buy Boxes won by Amazon and its special merchants are cases where Amazon did not offer the corresponding product at the lowest price. In the European marketplaces, these numbers are slightly lower (12.45% for France and 13.91% for Germany).
@@
Determining consumer preferences in digital markets in a reliable and reproducible manner is a challenging task. In particular, the data available to us from our empirical investigation of Amazon is insufficient to answer the question whether the Buy Box algorithm’s results correspond to consumer preferences. We therefore turn to a different empirical method that allows us to gain insights into consumers preferences: Using the Prolific crowdsourcing platform, we conducted a survey among 200 respondents (50 from each of the four countries).
@@
We continue our exploration of Amazon’s market practices by investigating its offer listing page, another choice architecture in the Amazon ecosystem. As described in the previous section, product pages on Amazon include a Buy Box that allows for quick product purchases and for which Amazon, if necessary, selects a seller from among the competing sellers offering a product. However, consumers are not bound to use the Buy Box. Amazon offers consumers a separate page which lists all the available offers for a product. Figure 3 displays an example of such an offer listing page. To get to the offer listing page, consumers can click a link that appears in a second rectangular box displayed below the Buy Box, see Figure 1. The offer listing page shows the offers from all the competing sellers in the increasing order of price and delivery charges. It also shows some seller performance metrics including its average user rating, the percentage of positive feedback, or the number of ratings.
@@
Second, we turn to a specific feature of Amazon’s approach to displaying customer ratings on the offer listings page that might be less innocuous, specifically because it might offer advantages to sellers who use its subsidiary fulfillment and/or delivery services (see Section 2). For these sellers, under certain circumstances, Amazon reserves the right to “strike through” negative feedback by consumers. For example, Amazon may strike through a seller feedback if “the entire comment relates explicitly to delivery experience for an order fulfilled by Amazon ” or “the entire comment is related to a delayed or undelivered order, which you shipped on time by using Buy Shipping” Amazon Seller Central (2022b). In both cases, in addition to the strike through, Amazon puts up a message taking responsibility for the inconvenience faced by the customer. Figure 4 shows examples of struck-through reviews. Amazon also states in its Fulfilled-by-Amazon and Amazon Shipment advertisements that negative reviews which have been struck through will not impact sellers’ performance metrics (Amazon Seller Central, 2022a, c). Importantly, this policy is unavailable for independent sellers that do not use Amazon’s fulfillment or delivery services. Hence, for the same customer inconvenience, sellers might be treated differently by Amazon depending on whether they purchase services from Amazon or not.
@@
We categorize the sellers in our dataset into three main groups: (1) Amazon Fulfilled, (2) Amazon Shipped, and (3) Merchant Fulfilled. The determination of seller membership in the first two groups was based on the text displayed when Amazon strikes through reviews. Sellers falling into the third group were identified if none of their reviews in the dataset were struck through. Subsequently, we conducted a manual inspection of their product pages to ascertain how they fulfill their orders.
We then assessed how often Amazon strikes through seller feedback pertaining to members of the first group, and how the striking through of feedback affected the relative ratings of Amazon Fulfilled sellers as opposed to Merchant Fulfilled sellers. Notably, more than half of the reviews in which Amazon Fulfilled sellers received only three stars or less was struck through in every country we investigated. For Amazon Shipped sellers, the numbers hover between 22% and 4% dependent on the country. Reviews for Merchant Fulfilled Sellers are never struck through.
The exclusion of struck-through feedback markedly changes how different types of sellers perform in ratings. For example, when excluding struck-through reviews, the overall ratings were best for Amazon Fulfilled sellers. When struck-through reviews are included in the evaluation, Amazon Fulfilled appear to have lower rankings than Merchant fulfilled sellers in India, the US, and Germany.
Consistent with our previous findings, these findings suggest that Amazon’s choice architecture favors sellers affiliated with Amazon. Different from the inclusion of the number of ratings as a performance metric, it seems less likely that the decision to exclude certain disfavorable ratings for Amazon-affiliated sellers from the calculation of seller performance is aligned with consumer preferences. At the very least, it results in the display of ratings that are not comparable across different types of sellers, thereby reducing their informativeness to consumers. At worst, it could distort consumers’ purchase decisions.
@@
% of 1st Preference Votes Amazon Metrics Rectified Metrics

### v2

@@ Introduction
From the perspective of individual users, digital platforms serve as both gateways to information and architects of their interaction spaces. Digital platforms wield considerable influence over what is commonly termed choice architectures (Thaler and Sunstein, 2008), the environments in which users navigate decision-making processes. Extensive research in psychology and related fields has demonstrated that the ability to manipulate choice architectures can profoundly shape user behavior and outcomes (Thaler and Sunstein, 2008). Digital platforms may use technical choice architectures expressed in algorithms to shape consumer preferences, search behavior and purchase patterns.
@@
That said, obtaining information about business models is usually not trivial. As a first step, a regulator or researcher seeking to assess the operations of a digital platform with the help of digital tools would typically try to access the platform through a regular web interface or engage in automated scraping of the platform’s content. However, interfaces of digital platforms have become more interactive, oftentimes rendering simple web access insufficient to capture the inner workings of a digital platform. Digital platforms and markets enabled by these platforms are growing more complex, making it harder to gain a proper understanding of platforms’ business practices.
@@
In response, empirically oriented computer science research has developed various approaches to audit digital platforms over the last few years. Sandvig et al. (2014) distinguish between “code audit[s]” (where researchers directly inspect a copy of the relevant algorithm), “noninvasive user audit[s]” (where users interacting with the platform answer questions about what they did online), “scraping audits” (where researchers query a platform repeatedly, observing results), “sock puppet audit[s]” (where researchers use computer programs to impersonate users), and “crowdsourced audit[s]” (where researchers employ users as testers). This development points towards a future of algorithmic auditing, where potential adverse effects of algorithmic systems are investigated and evaluated using algorithmic auditing tools (Sandvig et al., 2014; Bandy, 2021; Metaxa et al., 2021). “The Answer to the Machine is in the Machine", as Charles Clark pronounced in the 1990’s (Clark, 1999): As digital platforms are increasingly relying of algorithmic systems to govern their digital platforms, regulators will have to resort increasingly to algorithmic systems to audit compliance on these platforms.
@@
This article explores some of the opportunities and challenges of algorithmic enforcement, using Amazon as an example. Recently, Amazon’s business policies and deployed algorithmic systems have come under intense scrutiny due to Amazon’s vertical integration, or in other words, its involvement in multiple stages of the production and distribution process for its products and services. In this article, we present a series of empirical investigations into the extent to which Amazon engages in practices that are typically described as preferential treatment or self-preferencing. In these investigations, we analyze different features of various choice architectures that Amazon presents to its customers. We investigate how the service decides the offer that is featured in its Buy Box (now called “Featured Offer”); the design feature of its page listing various offers for the same product; how it selects products offered to customers that use the conversational product search technology available through its virtual assistant technology Alexa; and how it displays results in its related-item recommendation system.
@@
Over the last years, Amazon has emerged as the largest online e-commerce platform by connecting millions of sellers across the world to billions of customers (Saqib, 2023). On the Amazon platform, the different stakeholders can be primarily classified into three categories: (a) sellers on the platform they rely on Amazon to sell their products to consumers; (b) customers of the platform they rely on Amazon for fulfilling their purchase needs; and (c) the platform organization itself it mediates the interaction between the sellers and costumer.
@@
Second, the seller featured in the Amazon Buy Box and all other sellers are listed on a separate offer listing page (Dash et al., 2024). This page includes seller performance metrics such as average user ratings, percentage of positive feedback, and the number of ratings. This setup enables consumers to compare the sellers from whom they would like to purchase the product. We investigate the choice architecture of these offer listing pages in Section 4.2.
@@
While Amazon uses these and other choice architectures to design its platform, Amazon’s role is not restricted to a platform designer. In addition, Amazon sells directly to consumers in most of its marketplaces – for example, in the United States, France, or Germany – making it one of the major sellers on its own marketplace. This puts Amazon in direct competition with independent sellers for many products that are sold on Amazon. In some countries – for example, India – Amazon is not allowed to sell products on its marketplace directly. As a response, Amazon created joint ventures in India in collaboration with other retailers to sell products on Amazon.in. Cloudtail India and Appario Retail Private Ltd., for example, were two joint ventures which emerged as two of the biggest sellers on Amazon India. [^1] While Cloudtail ceased to offer products on Amazon in 2022, Appario remains in the business of selling products on Amazon. Both companies were sellers on Amazon India while the data presented in this article was collected.
@@
The intricate relationship between Amazon and sellers on Amazon Marketplace also plays out at the product level. Amazon manufactures its own private label products and sells them under different brand names (for example, AmazonBasics, Presto, or Solimo) (Dash et al., 2021). Note that only Amazon sells Amazon’s private label products on Amazon.
Finally, Amazon also maintains an advertising platform. For example, when a customer looks for a product on Amazon, the platform presents both organic and sponsored recommendations. While organic recommendations refer to the results that the algorithmic system produces in the desired context based on previous customer behavior, sponsored recommendations are sponsored product advertisements which are the result of second-price auctions in which sellers pay Amazon so that their product is displayed in a favorable position. Notably, Amazon also displays its own products among sponsored search results (Dash et al., 2021).
@@
The complex interactions between multiple players on Amazon are a result of continuous product development and innovation by Amazon over the last three decades. Originally, Amazon emerged as a book seller, entering the business of providing a platform for other sellers only later. Similarly, major online travel agencies such as Booking.com or Expedia over time became vertically integrated with meta-search platforms, enabling these travel agencies to compete with independent competitors on these platforms (Cure et al., 2022).
@@
When a digital platform such as Amazon sells its own products and products from affiliated producers on its online platform, this raises issues of vertical integration. Can a vertically integrated platform that operates both as platform provider and seller sufficiently distinguish between both roles? Or is there a risk that it has incentives to bias search results and recommendations in favor of its own or affiliated products, at the expense of products from competing independent sellers?
@@
Problems of vertical integration are not unique to digital platforms. The practice of supermarkets providing preferential shelf space to their own or affiliated products (see Tagliavini 2023, 202), and financial advisers earning commissions that might bias their investment recommendations (Cookson et al., 2021; Egan, 2019), both raise similar questions. From an economic perspective, vertical integration can have ambiguous effects on social welfare. Vertical integration of digital platforms may increase consumer search costs, discourage product innovation and prompt sellers to leave the market, as preferential treatment by the platform provider for own or affiliated products may make it more difficult for independent sellers to compete (see Farronato, Fradkin, and MacKay 2023). However, it is not clear that such welfare-decreasing effects will always dominate. By engaging in preferential treatment, a dominant vertically integrated platform provider may have incentives to invest in product quality, product design and innovation. It may also have an incentive to decrease prices and increase output because it can capture a larger fraction of the benefits from such actions (often referred to as the “internalization of double mark-ups,” see European Commission 2008, 7). Therefore, whether vertically integrated platform providers are beneficial or detrimental to economic welfare is a complex question that is often impossible to answer in the abstract, based solely on theoretical considerations (see, for example, de Cornière and Taylor 2019; Hagiu, Teh, and Wright 2022; Anderson and Bedre-Defolie 2022).
@@
These insights from industrial organization research have had an influence on competition law’s treatment of vertical integration. With the increasing influence of the Chicago School in antitrust, many jurisdictions have become reluctant to condemn cases of vertical integration as such. Rather, antitrust authorities investigate potential violations of antitrust laws in vertical relationships as part of a rule-of-reason analysis. Under rule-of-reason analysis, antitrust agencies and courts examine both the positive and negative effects of a certain conduct on competition in order to determine whether the conduct violates antitrust laws. For example, U.S. courts have shifted from a per se prohibition of resale price maintenance [^2] Compare Dr. Miles Medical Co. v. John D. Park & Sons Co., 220 U.S. 373 (1911) (treating a vertical price fixing scheme as per se illegal) with Leegin Creative Leather Prods., Inc. v. PSKS, Inc., 551 U.S. 877 (2007) (holding that all vertical distribution restraints – non-price and price – should be analyzed under the rule of reason). and territorial restrictions [^3] Compare United States v. Arnold, Schwinn & Co., 388 U.S. 365 (1967) (holding that vertical restraints limiting distributors to exclusive sales areas is per se illegal) with Cont’l T.V., Inc. v. GTE Sylvania, Inc., 433 U.S. 36 (1977) (applying a rule of reason analysis to non-price distribution restraints protecting interbrand competition). to a rule-of-reason analysis that focuses on the economic effects of vertical restraints. Relatedly, self-preferencing by a dominant firm is not a per se violation of European competition law (Art. 102 of the Treaty on the Functioning of the European Union, TFEU), but subjects to an effects test (Crémer, Montjoye, and Schweitzer, 2019, 7, 66).
@@
This effects-based approach to assessing vertical integration under antitrust laws has influenced the way authorities have addressed alleged preferential treatment by digital platforms. The General Court of the European’s recent Google Shopping decision provides an example. In this decision, the court reviewed the European Commission’s finding that Google had promoted its own comparison shopping service on its general search engine at the expense of rival services (see General Court, Case T-612/17, Nov. 10, 2021, ECLI:EU:T:2021:763). In its decision, the court pointed out that not all foreclosure is detrimental to competition, and that the mere extension of a dominant position to a neighbouring market is not necessarily anti-competitive (General Court, Case T-612/17, Nov. 10, 2021, ECLI:EU:T:2021:763, para. 157, 162). While the court ultimately confirmed the European Commission’s decision, the case demonstrates that whether self-preferencing violates European competition law can only be decided on a case-by-case basis. This analysis involves various different competition doctrines such as duty to deal, tying, margin squeeze, general abuse, or discrimination (see Colangelo 2023).
@@
Along these lines, the European Union recently introduced a regulatory system for digital markets that is based on an ex-ante approach where efficiency defenses raised by firms play a very limited role. With the Digital Markets Act (DMA), which came into full force in 2024, the European Union aims to make key digital markets fairer and more contestable by preventing large companies from abusing their market power and lowering barriers to the entry and expansion of smaller players. The DMA targets large digital platforms – so-called “gatekeepers” – such as Alphabet, Amazon, Apple, Meta, or Microsoft and imposes various obligations on them.
@@
Article 6(2) DMA prohibits digital platforms from using, in competition with other sellers, any non-public data that is generated or provided by those sellers. This prohibition is intended to countervail the ability of vertically integrated digital platforms to observe sales data from independent sellers with whom the platforms compete. The DMA also includes, among other provisions, prohibitions on combining data from different services belonging to the same company, protections for advertisers and publishers using the gatekeeper against monopolization trends, and provisions for ensuring interoperability and data access for business and individual users of platforms. If gatekeepers fail to abide by these obligations, the European Commission may impose fines of up to 10 % of the gatekeepers’ annual worldwide turnover (Art. 30(1) DMA). The Digital Market Act empowers the Commission to monitor the effective compliance of digital platforms with the obligations of the DMA, potentially using independent external experts, auditors and member state authorities (Art. 26 DMA). Article 21 DMA entitles the European Commission to require access to data and algorithms of digital platforms.
Going beyond the Digital Markets Act, the European Union also enacted the Digital Services Act (DSA), which came into effect in 2024 as well. This regulation provides a comprehensive framework regarding illegal content, transparent advertising, and disinformation of online platform providers. Article 26 of the DSA, for example, requires online platforms to be transparent about the nature and origin of advertisements. This obligation also applies to “sponsored links” displayed by digital platforms. Articles 25 and 27 of the DSA include rules about online interface design and recommender system transparency. The European Commission may impose fines of up to 6 % of the annual worldwide turnover of very large online providers (Art. 74(1) DSA), which includes Amazon, Alibaba, Apple, Google Shopping, and Zalando, among other providers. [^4] For a list of very large online platforms as designated by the European Commission, see https://digital-strategy.ec.europa.eu/en/policies/list-designated-vlops-and-vloses.accessed April 02, 2024
@@
In the following section, we summarize a series of empirical investigations into various of Amazon’s choice architectures undertaken by several co-authors of this article (Dash et al., 2021, 2022, 2024). The purpose of this section is to describe their main findings and distill important insights about the promises and challenges of algorithmic enforcement of platforms.
@@
We start our exploration of Amazon’s market practices by investigating how the service decides which offer to feature in its Buy Box [^5] Summarized from Dash et al. (2024). (now called “Featured Offer”). A typical Amazon product page has a rectangular box on the top-right corner, right next to the title of the product. This box is popularly known as the Buy Box. Figure 1 shows an example of such a Buy Box. The box contains the price of the product, a delivery estimate, and two buttons that read: “Add to Cart” and “Buy Now”. The box also lists the name of a seller. In the Buy Box depicted in Figure 1, the seller is Appario Retail Private Ltd.
@@
Given the importance of the Buy Box for transactions on Amazon’s platform, Amazon exerts considerable power over customers’ purchase decisions through its algorithm that selects the seller that appears in the Buy Box. This decision is trivial if there is only one seller that offers a given product. However, in practice, multiple sellers can and often do offer the same product. In this scenario, Amazon selects the seller to be featured in the Buy Box through a proprietary algorithm known as the Featured Offer Selection algorithm, also commonly known as the Buy Box algorithm. The precise structure of this algorithm is unknown, making it a subject of speculation (and concern) among sellers on Amazon (Kaneshiro, 2020). Amazon presents some clues and recommendations on its seller interface regarding what factors its algorithms take into account in determining the winner of the Buy Box. These factors include, for example, competitive pricing and seller performance metrics.
@@
To investigate this question, we first compare the price offered by Amazon and rival sellers in cases in which Amazon won the Buy Box. We find considerable numbers of instances in which Amazon won the Buy Box despite offering the product at a price that is higher than that offered by at least one rival seller: In the US and in India, 25% of the Buy Boxes won by Amazon and its special merchants are cases where Amazon did not offer the corresponding product at the lowest price. In the European marketplaces, these numbers were slightly lower (12.45% for France and 13.91% for Germany).
@@
Determining consumer preferences in digital markets in a reliable and reproducible manner is a challenging task. The data available to us from our empirical investigation of Amazon is insufficient to answer the question whether the Buy Box algorithm’s results correspond to consumer preferences. We therefore turn to a different empirical method that allows us to gain insights into consumers preferences: Using the Prolific crowdsourcing platform, we conducted a survey among 200 respondents (50 from each of the four countries).
@@
We continue our exploration of Amazon’s market practices by investigating its offer listing page, another choice architecture in the Amazon ecosystem [^7] Summarized from Dash et al. (2024).. As described in the previous section, product pages on Amazon include a Buy Box that allows for quick product purchases and for which Amazon, if necessary, selects a seller from among the competing sellers offering a product. However, consumers are not bound to use the Buy Box. Amazon offers consumers a separate page which lists all the available offers for a product. Figure 3 displays an example of such an offer listing page. To get to the offer listing page, consumers can click a link that appears in a second rectangular box displayed below the Buy Box, see Figure 1. The offer listing page shows the offers from all the competing sellers ordered by price and delivery charges. It also shows some seller performance metrics including its average user rating, the percentage of positive feedback, and the number of ratings.
@@
Second, we turn to a specific feature of Amazon’s approach to displaying customer ratings on the offer listings page that might be less innocuous, specifically because it might offer advantages to sellers who use its subsidiary fulfillment and/or delivery services (see Section 2). For these sellers, under certain circumstances, Amazon reserves the right to “strike through” negative feedback by consumers. For example, Amazon may strike through a seller feedback if “the entire comment relates explicitly to delivery experience for an order fulfilled by Amazon ” or “the entire comment is related to a delayed or undelivered order, which you shipped on time by using Buy Shipping” Amazon Seller Central (2022a). In both cases, in addition to the strike through, Amazon puts up a message taking responsibility for the inconvenience faced by the customer. Figure 4 shows examples of struck-through reviews. Amazon states in its Fulfilled-by-Amazon and Amazon Shipment advertisements that negative reviews which have been struck through will not impact sellers’ performance metrics (Amazon Seller Central, 2022b, c). Importantly, this policy is unavailable for independent sellers that do not use Amazon’s fulfillment or delivery services. Hence, for the same customer inconvenience, sellers might be treated differently by Amazon depending on whether they purchase services from Amazon or not.
@@
We categorized the sellers in our dataset into three main groups: (1) Amazon Fulfilled, (2) Amazon Shipped, and (3) Merchant Fulfilled. The determination of seller membership in the first two groups was based on the text displayed when Amazon strikes through reviews. Sellers falling into the third group were identified if none of their reviews in the dataset were struck through. Subsequently, we conducted a manual inspection of their product pages to ascertain how they fulfill their orders.
We then assessed how often Amazon struck through seller feedback pertaining to members of the first group, and how the striking through of feedback affected the relative ratings of Amazon Fulfilled sellers as opposed to Merchant Fulfilled sellers. Notably, more than half of the reviews in which Amazon Fulfilled sellers received only three stars or less was struck through in every country we investigated. For Amazon Shipped sellers, the numbers hover between 22% and 4% dependent on the country. Reviews for Merchant Fulfilled Sellers were never struck through.
The exclusion of struck-through feedback markedly changes how different types of sellers perform in ratings. For example, when excluding struck-through reviews, the overall ratings are best for Amazon Fulfilled sellers. When struck-through reviews are included in the evaluation, Amazon Fulfilled appear to have lower rankings than Merchant fulfilled sellers in India, the US, and Germany.
Consistent with our previous investigations, these findings suggest that Amazon’s choice architecture favors sellers affiliated with Amazon. Different from the inclusion of the number of ratings as a performance metric, it seems less likely that the decision to exclude certain disfavorable ratings for Amazon-affiliated sellers from the calculation of seller performance is aligned with consumer preferences. At the very least, it results in the display of ratings that are not comparable across different types of sellers, thereby reducing their informativeness to consumers. At worst, it could distort consumers’ purchase decisions.
@@
Percentage of 1st Preference Votes Amazon Metrics Rectified Metrics